OpenStudy (anonymous):

Two chords intersect inside a circle. The lengths of the segments of one chord are 4 and 6. The lengths of the segments of the other chord are 3 and

Directrix (directrix):

4*6 = 3*x 3x = 24 x = ?

Directrix (directrix):